using the five-step problem-solving process
A rancher outside Houston decides to build a silo inside his corral. The rectangular corral measures 23 m by 17 m and the silo's base is circular with a diameter of x meters. Find a polynomial for the remaining area of the corral (in square meters) after the silo is built.

Answer by mananth(16946)

Length of corral = 23 m
width of corral = 17m
..
Area of Corral = 17*23 = 391 m^2
..
diameter of silo = x
radius = x/2
..
Area of circle = pi*r^2
Area = pi*(x/2)^2
..
Area of corral - area of silo = area of remaining area
391- pi*(x/2)^2